OpenStudy (mathmath333):

A lion takes 5 leaps for every 4 leaps of tiger but 3 leaps of tiger are equal to 4 leaps of lion. What is ratio between speed of lion to tiger?

OpenStudy (anonymous):

I can be wrong :D x lion time :5x=4y distance:4x=3y. dividing both. 4/5 V(lion)=3/4

OpenStudy (rational):

15 : 16 looks right

OpenStudy (mathmath333):

lol how did u get that @rational

OpenStudy (anonymous):

Yea it is the same answer ,but much faster :D

OpenStudy (rational):

4/5 V(lion)=3/4 V(tiger) V(lion)/V(tiger) = 15/16 ?

OpenStudy (rational):

just used Catch.me's work haha!

OpenStudy (mathmath333):

lol i didnt see at first

OpenStudy (rational):

a less mathematical way of working it : `A lion takes 5 leaps for every 4 leaps of tiger` is same as `A lion takes 15 leaps for every 12 leaps of tiger`

OpenStudy (rational):

since we're given that `3 leaps of tiger are equal to 4 leaps of lion`, 12 leaps of tiger would equal 16 leaps of lion

OpenStudy (mathmath333):

haha that was too confusing for me

OpenStudy (rational):

so A lion takes `15 lion leaps` for every `16 lion leaps` of tiger confusingly the ratio is 15:16 :P

OpenStudy (mathmath333):

books gives \(15:16\) as correct

OpenStudy (rational):

yaaay!